### v1

# The Popularity Hypothesis in Software Security: A Large-Scale Replication with PHP Packages

Abstract
There has been a long-standing hypothesis that a software’s popularity is related to its security or insecurity in both research and popular discourse. There are also a few empirical studies that have examined the hypothesis, either explicitly or implicitly. The present work continues with and contributes to this research with a replication-motivated large-scale analysis of software written in the PHP programming language. The dataset examined contains nearly four hundred thousand open source software packages written in PHP. According to the results based on reported security vulnerabilities, the hypothesis does holds; packages having been affected by vulnerabilities over their release histories are generally more popular than packages without having been affected by a single vulnerability. With this replication results, the paper contributes to the efforts to strengthen the empirical knowledge base in cyber and software security.
keywords:
Software security, software vulnerability, replication, web applications, software ecosystems, dependencies, downloads, installs, stars, forks
1 Introduction
The paper examines what can be called a popularity hypothesis in empirical software security research; an assumption that a software’s popularity affects its security or insecurity. The hypothesis has been previously examined with Java packages, the conclusion being that popularity is not a reliable indicator of software security [1]. The present paper revisits this hypothesis with a large-scale analysis of open source software packages written in the PHP programming language. Given the existing work, the paper is by definition a replication.
Within the domain of replication studies, the paper can be classified as a dependent or a conceptual replication: although both the data and methods are different, the paper is designed with an explicit reference to previous work in mind [2, 3, 4]. The methodological difference can be justified on the grounds that the replicated study was based only on correlation analysis [1]. In contrast, the paper uses also regression analysis and classification algorithms, which can be argued to yield stronger empirical evidence in general. The recently discussed generalizability problems in empirical software engineering [5] justify the use of a different dataset. If the hypothesis holds, or does not hold, with two distinct datasets, there is either stronger or weaker evidence for it, respectively.
The underlying replication logic focuses on the previous work’s conclusion—it is important to know whether or not the hypothesis holds with a different dataset and different methods because the hypothesis is rather fundamental in many ways. It also frequently appears in popular discourse. The primary historical example would be Microsoft Windows; not so long ago, it was commonly discussed and speculated by both laypeople and professionals that the popularity of the operating system was a definite factor behinds its historical insecurity. To quote from a famous security engineering textbook, most “malware writers targeted Windows rather than Mac or Linux through the 2000s and 2010s as there are simply more Windows machines to infect”, and, furthermore, the “model replicated itself when smartphones took over the world” [6, p. 296]. Although there is no particular reason to question these assertions, it is still worth remarking that they are not actually backed by any empirical evidence in the textbook. This point motivates a need to examine the hypothesis further empirically.
Given a distinction between “hard” and “soft” theories in empirical software engineering [7], the popularity hypothesis is on the side of hard theories because it presumes a causal relation between a cause (popularity) and effect (insecurity), but the theorizing involved is still rather soft. In other words, assuming that the hypothesis holds, the underlying explanations remain open to a debate.
The starting point for an explanation given by what is now known as (information or cyber) security economics would be incentives [6]. That is, cyber criminals and other adversaries have an incentive to target popular software products, packages, and projects, whether the targeting is about finding vulnerabilities from those or explicitly attacking them. Then: even though both commercial software vendors and open source software projects have an incentive to develop secure code and provide defenses, it is frequently discussed in the literature that software security and cyber security in general are not prioritized because there are no pressing incentives to do so [8]. A slightly different theorization path involves the famous so-called Linus law; that all bugs are shallow when there are many eyes involved to look at code [9]. From this alternative perspective, there are also empirical studies indicating that widely used and hence popular open source software packages often contain more reported bugs and vulnerabilities than less popular packages [10, 11, 12]. These empirical studies further reinforce the need and motivation for a replication of the popularity hypothesis study already mentioned earlier. The italics placed upon the word reported are also worth briefly explaining. The reason for the italics is that empirical observations cannot generally prove that something is secure, although these can be used to make claims about insecurity [13]. This fundamental point should be kept in mind throughout the paper; only reported security issues are observed. By implication, the popularity hypothesis in itself may not necessarily tell everything, or in some cases even anything, about software security per se.
The following three contributions are made:
  1. 1.
    The paper contributes to strengthening of the empirical knowledge base in cyber and software security by evaluating the robustness of existing knowledge.
  2. 2.
    The paper’s replication advances the use of good practices in empirical cyber and software security research.
  3. 3.
    With respect to the replicated study [1], the paper examines a substantially larger dataset and uses more sophisticated methods for its analysis.
The paper’s remainder is structured into a conventional format. The opening Section 2 further frames the paper with a larger scope of related work. Then, the empirical materials and the methods for examining these are elaborated in Section 3. This section also clarifies what is meant by software popularity; in general, it is operationalized as a degree to which a package is widely adopted and actively engaged within its ecosystem. The subsequent Section 4 presents the replication results after which conclusions follow in the final Section 5.
2 Related Work
The paper’s reference domain is empirical software engineering focusing on software security. The domain is large, but with some suitable framings, it is possible to narrow it down. The first obvious framing is toward empirical research. The second equally obvious framing is toward the PHP programming language. With these two framings, a few examples can be given about implicitly related work. For instance, technical debt of PHP packages for web applications has been investigated empirically [14]. Regarding web applications and websites in general, another example would be a large-scale empirical study investigating the adoption of PHP interpreter releases among website deployments [15]. A third framing can be done toward software security. The previous two examples are in the scope also with this further software security framing—technical debt is related to software quality, which is a superset for software security, while the adoption (or a lack thereof) of new PHP interpreter releases contains security risks in case many (or some important) websites do not update their interpreters for a reason or another. The reason for the risks is that also the PHP interpreter has seen many vulnerabilities over the years.
A lot of empirical research has also been done to examine the security of software written in PHP more generally. For instance, the security of plugins for popular PHP-based web frameworks has been investigated [16, 12]. Regarding software vulnerabilities and their underlying weaknesses, and despite the availability of defensive solutions [17], cross-site scripting and input validation more generally have been the most typical weaknesses in PHP software [12, 18]. Though, it can be noted that this observation is hardly unique to PHP. In other words, the same weaknesses typically lead the scoreboards also with other popular interpreted programming languages used in the web domain [19]. Having said that, it should be emphasized that empirical software security research on PHP applications is not limited only on vulnerabilities and weaknesses. A good example would be a forensics investigation of PHP applications in containerized cloud environments based on empirical log analysis [20]. Furthermore, a third framing can be done toward replication research.
Despite a long debate on a real, perceived, or alleged need to align with empirical sciences in cyber security research [13], the empirical foundations are arguably still rather limited. Like in empirical software engineering [21], which also partially frames the current paper, systematic literature reviews have improved the systematization of knowledge but actual cumulation of empirical evidence, whether done through meta-analyses or by other means, has been limited. Many reasons for this limitation could be pointed out and further speculated. Among other things: like in many fields, there has been a lack of incentives in cyber security research to share datasets [22]. Although there are replications also in cyber security [4], the data sharing limitation has supposedly contributed to their volume. To this end, the paper contributes to the efforts to strengthen the empirical knowledge base in cyber security.
A fourth and final clarifying framing can be done toward research on software ecosystems and particularly the security or insecurity of software distributed via them. The empirical security-oriented ecosystem research is again vast. A recurring theme is that software packages distributed in software ecosystems via language-specific package managers are generally of poor equality, often containing security issues of various kinds, including concrete vulnerabilities [23]. In addition, the ecosystems exhibit a risk of malware being uploaded, as is often done together with so-called typo-squatting [19, 24]. The underlying security risks are intensified by the heavy use of software dependencies in these ecosystems [25]. Given this background, it is understandable that there are also various ongoing funded projects to improve the ecosystems and the supply chain security of open source software in general [8]. The software ecosystem research and dependencies are also important to underline because the concept of (software) popularity is closely related to them [26, 27]. The framing toward software ecosystems is also relevant as it is also directly related to the dataset examined. On that note, the materials and methods should also be elaborated.
3 Materials and Methods
In what follows, the large-scale PHP dataset for the replication is first briefly elaborated. Afterwards, the computational methods for analyzing it are described.
3.1 Data
The dataset was collected on December 2024 from Packagist [28], a repository for PHP software packaged with the Composer package manager [29]. The dataset is available online for any interested party, possibly regarding further replication work [30]. In total, $n=381,993$ packages are covered in the dataset. The dataset is composed of eight simple meta-data metrics provided on the Packagist’s website: security, installs, dependents, suggesters, stars, watchers, forks, and releases. Of these metrics, all except security and releases can be interpreted to proxy slightly different dimensions of software popularity. The outlying releases metric counts the number of releases made for a given open source PHP package, as reported on the Packagist’s website. All metrics have a continuous scale.
While installation amounts, installs, provide a relatively straightforward and widely used metric for software popularity [11, 26, 27, 12], the dependents metric conveys a package’s popularity in terms of incoming edges in a dependency network between packages [26]. With this metric a package is seen as popular when many other packages depend on it. Both installation (or download) counts and dependencies have also been used to help security-related decision-making about funding applications for open source software projects [8]. Although documentation is unfortunately lacking, the suggesters metric generally proxies “soft” dependencies; for each package, Packagist recommends also further packages in addition to the “hard” dependencies, as captured by dependents, without which a package does not install or function properly.
The stars, watchers, and forks metrics are explicitly tied to functionality provided by GitHub on which almost all of the actual development of the PHP packages occurs today. In general, with “stars” developers can flag packages and projects they perceive as interesting or relevant. By adding oneself as a “watcher” to a project, one can get information about changes in the project. The forks metric, in turn, refers to the number of times a given project was forked, as has been common in the open source world throughout the decades [31]. According to practitioner surveys, all three metrics have been seen as relevant for proxying popularity [32]. Finally, the dependent metric is security, which counts and records vulnerabilities reported for a package. Many (but not all) of the vulnerabilities tracked on Packagist are identified with Common Exposures and Vulnerabilities (CVEs), while the associated security advisories typically point toward GitHub. In this regard, it can be noted that no additional data was retrieved to validate the CVE-referenced vulnerability information provided on Packagist. Nor was GitHub queried to retrieve further software development data.
3.2 Methods
The primary methodology is based on binary classification. In addition, a few descriptive statistics and a brief ordinary least squares (OLS) regression analysis are presented. As these do not require a particular exposition, the following discussion concentrates on the classification.
3.2.1 Classification
Imbalance is a highly typical problem in computational cyber security applications [33, 34]. Although not a problem for the current paper, not only is cyber security data often imbalanced but it also often contains missing values [35]. Even without a problem of missing values, as will be soon seen in Section 4, the dataset examined can be characterized to contain extreme imbalance. In other words, only a small minority of the PHP packages examined has been affected by one or more vulnerabilities.
To tackle the issue, four computational solutions are used. The first is oversampling or upsampling, as it is also known. It balances a training set by replicating the minority class, which in the present context refers to packages that have been affected by one or more vulnerabilities during their entire release histories. The second is undersampling or downsampling; it eliminates observations from the majority class (non-vulnerable packages) until a training set is balanced. Although there are variants that use synthetic data [33], both the oversampling and undersampling solutions are computed with random samples.
Then, the third and fourth solutions are based on the SMOTE [36, 37] and ROSE [38, 39] algorithms. The former uses synthetic data to mimic observations in the minority class. The $k$-nearest neighbors algorithm is used to interpolate the synthetic observations. The latter uses similar ideas with bootstrapping methods. In general, particularly the SMOTE algorithm is commonly seen as a today’s de facto solution for handling imbalanced data [40]. Although numerous competing algorithms have been developed over the years, SMOTE therefore suffices as a good baseline for the classification computations. Furthermore, it should be emphasized that the goal is not about seeking maximum performance but about testing of a hypothesis.
In the same vein, actual classification is done with three conventional algorithms: naïve Bayes [41], (boosted) logistic regression [42], and random forest [43]. All are well-known and thus require no particular elaboration. Instead, the implementations used are worth remarking; these are all R packages implemented to work with the caret package [44]. Default parameters were used for all algorithms. The only exception is the random forest classifier implementation for which the maximum number of trees had to be restricted to one hundred due to memory constraints.
3.2.2 Performance Metrics
Imbalance affects also metrics for evaluating classification performance. Among other things [39], the traditional accuracy metric is biased and misleading in this context. Although the $F_{1}$ metric is often used as an alternative, it too can be problematic in cyber security applications [45]. Therefore, three alternative performance metrics are used.
First, a simple balanced accuracy (BA) has been used in existing work [46]. It is given by
$\textmd{BA}=(\textmd{TPR}+\textmd{TNR})~{}/~{}2,$ (1)
where the true positive rate (TPR) and the true negative rate (TNR) are defined as
TPR $=\textmd{TP}~{}/~{}(\textmd{TP}+\textmd{FN})~{}~{}~{}~{}~{}\textmd{and}~{}$ (2)
TNR $=\textmd{TN}~{}/~{}(\textmd{TN}+\textmd{FP}),$ (3)
where, in turn, TP refers to true positives, FN to false negatives, TN to true negatives, and FP to false positives. The TPR metric is also known as a recall or a sensitivity and the TNR metric as a specificity.
Second, a G-mean metric has also been used in the imbalanced classification context [47, 48]. It is defined via the conventional metrics in (2) as
[…]
The BA and G-mean metrics both vary in the unit interval. Higher values are better.
Third, the conventional mean square error (MSE), which is sometimes also known as the Brier score [47], has further been used in the imbalanced classification context. It is given by the usual formula:
$\textmd{MSE}=\frac{1}{n}\sum^{n}_{i=1}(\widehat{p}_{i}-p_{i})^{2},$ (5)
where, in the present context, $\widehat{p}_{i}$ is an estimated probability for a package’s release history having been affected by at least one vulnerability, while $p_{i}$ is the actual probability; $p_{i}=1$ in case the $i$:th package has been affected by at least one vulnerability. Unlike with BA and G-mean, lower values are better.
3.2.3 Computation
Three models are computed for each classifier and each of the four balancing solutions. Thus, in total, $3\times 3\times 4=36$ computations are carried out. Evaluation is done according to the three performance metrics.
The first model contains only releases. As was discussed in Subsection 3.1, it is the only explanatory metric that is not explicitly related to popularity of open source software. Therefore, the first model serves as a baseline for comparing the other two models. If the hypothesis holds, adding the popularity metrics should increase performance.
The second model contains two explanatory metrics: again releases and a sum variable based on the arithmetic mean of the six popularity metrics. Cronbach’s standardized $\alpha$-coefficient [49] is $0.64$ for this sum variable. Given the maximum of one, the value is not particularly high but still sufficient for the present purposes. Adjectives such as acceptable, satisfactory, and sufficient have been used to describe comparable values in existing research [50]. The third full or unrestricted model contains all available information in the dataset. In other words, the seven metrics described earlier in Subsection 3.1 are used in this full model. Alongside the hypothesis, the prior expectation is that the model’s performance is better compared to the barebone first model.
In addition, for a reason soon discussed, a $\ln(x+1)$ transformation is applied to all metrics, including the security metric that is truncated into a binary-valued metric later on for the classifications. Regarding other computational details, a $10$-fold cross-validation is used. Testing is done with a random sample containing 25% of the dataset. Due to the logarithm transformation, neither scaling nor centering are used for the explanatory metrics. With these notes in mind, the results can be disseminated next.
4 Results
The presentation of the results is done in three steps. A few descriptive statistics are first presented. Then, the hypothesis is briefly further probed with regression analysis. The classification results end the presentation.
4.1 Descriptive Statistics
The presentation of the results can be started by reiterating the methodological points. Thus, the extreme imbalance is best illustrated with plain numbers: of the over $381$ thousand PHP packages, only $777$, or about $0.2\%$, have been affected by one or more vulnerabilities over their release histories. When taking a look at Table 1, it can be further concluded that the majority (57%) of the $777$ packages with vulnerabilities has been affected by a single vulnerability. As has also been previously observed with PHP packages [12], the tail is relatively long; there are outlying packages that have been affected by nine or more vulnerabilities. A similar observation applies to the remaining metrics; they all have long tails. Thus, the $\ln(x+1)$ transformation is generally justified for analyzing the dataset.
Table 1: Vulnerabilities Across the PHP Packages
Number of vulnerabilities
0 1 2 3 4 5 6 7 8 $\geq$ 9
Frequency $381,216$ $433$ $114$ $51$ $27$ $23$ $15$ $7$ $14$ $93$
Share (%) $99.797$ $0.113$ $0.030$ $0.013$ $0.007$ $0.006$ $0.004$ $0.002$ $0.004$ $0.024$
Figure 1: Correlations in the Dataset
Figure 2: Means Across Vulnerability Groups (Welch’s [51] approximation for unequal variances; all $t$-tests statistically significant at $p<0.001$)
Figure 3: Regression Results (OLS, full sample)
Regarding multicollinearity, Fig. 1 displays (Pearson’s) correlation coefficients across all metrics in the dataset by using the logarithm transformation. As can be seen, all coefficients have positive signs and some of these are relatively large in their magnitudes. These correlations are nothing surprising as such. Particularly when operating with large behavioral datasets—a domain to which the dataset examined can also be seen to belong, everything tends to be correlated with everything [52]. In any case, stars, watchers, and installs are moderately or even strongly correlated with each other on one hand and installs and releases on the other, to use the thresholds and adjectives for these from the replicated study [1]. Of these correlations, particularly the latter two seem sensible in a sense that more releases may cause more installs. These correlations may affect the classifications because particularly the naïve Bayes and logistic regression rely on an independence assumption. Though, even naïve Bayes seems to still perform well under multicollinearity [53]. Therefore, it is more relevant to continue by pointing out the only modest magnitudes between security and rest of the metrics. This observation might be taken as a prior expectation that the hypothesis may not hold.
However, the descriptive results in Fig. 2 tell a different story. These are based on conventional $t$-test estimates using a correction for unequal variances, which are evident also according to the Leneve’s classical test [54]. As can be seen, all means are different in the two groups separated by the truncated, binary-valued security metric. In other words, both popular packages and packages with long release histories seem to have been affected by vulnerabilities over the years. This observation provides a good motivation to briefly also report regression analysis results before continuing to the classification results.
4.2 Regression Analysis
Regression results can shed a little more light on the earlier correlation results in Fig. 1. Thus, coefficients from three OLS regressions for the three models noted in Subsection 3.2.3 are shown in Fig. 3. The logarithm transformation is again used for all metrics. All coefficients are statistically significant at a 99% confidence level, which is hardly surprising due to the sample size [55]. As could be furthermore expected, the performance increases the more there are metrics. The full model yields an adjusted $R^{2}=0.045$, meaning that roughly about five percent of the total variance is explained by the seven metrics. When keeping in mind that there are only seven hundred seventy seven observations for which security attains a value larger than zero, the performance is not necessarily that bad. Of the coefficients for the individual metrics, the one for suggesters stands out. Interestingly, the coefficient for installs has a negative sign, which seems to contradict the earlier results in Fig. 2. All in all, nevertheless, the evidence is again on the side of supporting the hypothesis. The magnitudes and signs of the coefficients are presumably linked to the correlations between the explanatory metrics.
4.3 Classification
The classification results are summarized in Tables 2, 3, and 4 for the three classifiers. In each table, the second, third, and fourth columns denote the three models estimated. To recall: the first model is the barebone one with just the releases metric, the second model adds the sum variable of the popularity metrics, and the third model uses all seven metrics individually. Then, the rows are arranged according to the three performance metrics described in Subsection 3.2.2. For each metric, the results are reported according to the four balancing solutions.
Table 2: Naïve Bayes Classification Results1
Model 1. Model 2. Model 3.
BA
Oversampling 0.771 0.726 0.749
Downsampling 0.771 0.736 0.803
[…]
ROSE 0.771 0.726 0.809
G-mean
Oversampling 0.759 0.678 0.747
Downsampling 0.759 0.693 0.803
[…]
ROSE 0.759 0.678 0.807
MSE
[…]
Downsampling 0.292 0.442 0.206
SMOTE 0.284 0.463 0.199
ROSE 0.278 0.460 0.219
  • 1.
    1 The best values are colored in each of the three rowwise panels.
Table 3: Boosted Logistic Regression Classification Results1
Model 1. Model 2. Model 3.
BA
Oversampling 0.731 0.754 0.805
Downsampling 0.793 0.633 0.761
SMOTE 0.731 0.759 0.816
[…]
G-mean
Oversampling 0.695 0.722 0.796
Downsampling 0.788 0.522 0.732
SMOTE 0.695 0.727 0.814
[…]
MSE
Oversampling 0.492 0.383 0.260
Downsampling 0.299 0.593 0.378
SMOTE 0.492 0.366 0.194
[…]
  • 1.
    1 The best values are colored in each of the three rowwise panels.
Table 4: Random Forest Classification Results1
Model 1. Model 2. Model 3.
BA
Oversampling 0.505 0.543 0.508
Downsampling 0.704 0.689 0.772
SMOTE 0.524 0.599 0.525
[…]
G-mean
Oversampling 0.484 0.319 0.125
Downsampling 0.694 0.622 0.746
SMOTE 0.405 0.525 0.231
[…]
MSE
[…]
Downsampling 0.327 0.496 0.314
SMOTE 0.145 0.094 0.008
ROSE 0.464 0.464 0.104
  • 1.
    1 The best values are colored in each of the three rowwise panels.
With these notes, the dissemination of the classification results can be started by noting that the balancing solutions improve the performance significantly. Although not reported, plain unbalanced classifications yield much worse performance, regardless of the three classifiers and the three models. Regarding the balancing solutions, it seems that the ROSE algorithm is better than the rest, although SMOTE does well with the Naïve Bayes classifier, and plain oversampling yields the best performance in one outlying case. As for the classifiers, the boosted logistic regression outperforms the Naïve Bayes and random forest classifiers according to the BA and G-mean metrics. According to these two performance metrics, the highest value reached is $0.840$, which can be interpreted as a decent value in the extreme imbalance context operated.
Most importantly, the best values colored in each table refer to the third model. In all cases these are much better than with the first model. The popularity hypothesis can be taken to hold also according to the classification results.
5 Conclusion
The paper revisited a popularity hypothesis in empirical software security research; an assumption that a software’s popularity can explain to some degree the software’s insecurity. According to the results presented, the hypothesis can be concluded to hold: popular PHP packages have been affected by vulnerabilities throughout their release histories more often than unpopular PHP packages. By implication, the paper also fails to replicate a previous study’s conclusion, although it should be noted that the study strictly speaking only concluded that popularity is not a reliable indicator of software security [1]. Although a decent classification performance was obtained even with the limited amount of explanatory information used, the results reported should not be taken to mean that popularity alone could explain the security or insecurity of packages written in the PHP programming language.
The confirmation of the hypothesis has implications also for empirical research more broadly. Recently, it has been argued that probabilistic random sampling should be preferred in empirical software engineering to fix a real or perceived generalizability crisis [5]. The results presented cast a doubt upon the argument. The unconditional probability of picking a non-vulnerable PHP package from the almost entire Packagist population observed is $0.998$. In other words, it would be very likely that only non-vulnerable packages would end up in a sample picked randomly. By implication, it is necessary to return to the concept of reported vulnerabilities discussed in the introduction. Thus, should one consequently believe that the $381,216$ packages in the sample without reported vulnerabilities are free of security issues and generally of high quality? Although no definite answers can be given, many would likely prefer a negative answer to the question. If this answer is accepted, the why-question subsequently emerges. As was noted, the famous Linus law probably has something to do also with reporting of vulnerabilities in large software ecosystems. If a package is particularly popular, it may be more likely that some developers actually also review and even audit its code base.
The confirmation of the popularity hypothesis has also other implications. As was noted, popularity metrics have also been used to grant cyber security funding for open source software projects. As the funding grants have involved also testing and security audits [8], it may be that the evidence for the popularity hypothesis also strengthens in the future in a sense that even more vulnerabilities are reported for popular open source software projects. The same point applies with respect to automated tools. For instance, also large-scale security scanning efforts of open source software projects have used popularity and related metrics to select preferable samples for scanning [56]. Audits and scanning are also important topics because they are arguably closer to “real” software security than what is available through observing reported vulnerabilities. To this end, it can be argued that also construct validity should be revisited, discussed, and debated in empirical software security research. In the meanwhile, it suffices to reiterate an argument that reported vulnerability counts should be approached with care when making practical choices about open source software packages [19]. This argument reinforces the earlier remark about long-standing philosophical and theoretical problems in applying principles of empirical sciences to cyber security research.
Finally, both the replicated study and this replication suffer from potential generalizability issues. In other words, it remains unclear whether the hypothesis would, or would not, hold with a further different dataset. As said, however, it is also generally unclear how the generalizability problem should be addressed—and whether it is even possible to address it in empirical software security research and empirical software engineering in general. Regardless, as it stands, there is now conflicting evidence about the popularity hypothesis. It is up to a reader to make the final verdict about whether software’s popularity affects its insecurity in light of the evidence put forward.